Part A A spring has an unstretched length of 15 cm. When an 85 g ball is hung from it, the length increases by 7.0 cm. Then the ball is pulled down another 7.0 cm and released. What is the spring constant of the spring? Express your answer in newtons per meter. N/m Submit Request Answer Part B What is the period of the oscillation? Express your answer in seconds. Submit Request Answer Provide Feedback
Answer #1

Given : M = 85 g = 0.085 kg , x = 7 cm = 0.07 m

Solution:

(a) Spring constant of the spring

We know that , F = -kx

Where F is the force due to gravity , k is the spring constant.

F = Mg = (0.085 kg)(9.81m/s2) = 0.834 N

i.e. 0.834 N= (k)(0.07 m)

k = (0.834 N)/(0.07 m) = 11.91 N/m

Answer : k = 11.91 N/m

(b) Period of oscillation

Period of oscillation is given by: T = 2\pi\sqrtM/k

T = 2(3.14)\sqrt(0.085/11.91)

= 0.53 s

Answer: T = 0.53 s